OpenStudy (anonymous):

how do you solve |2x-7|=11 I forgot the absolute value rules...

OpenStudy (anonymous):

THis splits into two equations. \[|2x - 7| = 11 \implies\]\[2x - 7 = 11 \space \space \space or \space \space \space 2x - 7 = -11\]

OpenStudy (anonymous):

so 9<x<9 are you serious?

OpenStudy (anonymous):

@niravshah08 This is an absolute value equation and not an absolute value inequality.

OpenStudy (anonymous):

x=9 and x=-2 ?

OpenStudy (anonymous):

so x is always positive that's the rule?

OpenStudy (anonymous):

Well, when you plug it in, you have to understand that whatever comes out of the absolute value bars will be negative. For instance, for this one: \[|2x - 7| = 11\]You have to get the left equal to 11 right? THerefore, one equation is 2x - 7. However, if the left side inside the absolute value were to equal -11, it would also be correct since |-11| = 11 right? That's why this happens.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

ohhhhhhhh yeah I remember it now thanks again, you're really helpful today!!

OpenStudy (anonymous):

np :) and glad to help!